# New GPU developments in the Madgraph CUDACPP plugin: kernel splitting, helicity streams, cuBLAS color sums

### v1

@@ 2.2 A closer look at the internals of the ME engine
    Arithmetic operations and other computational functions. Finally, the decoupling of data access from actual calculations has probably been one of the most important aspects of the whole CUDACPP software design. Just like the memory allocation classes, also the helicity amplitude functions like VXXXXX, VVV1P0_1 or FFV1_0 have almost [^1] One notable exception is that some assumptions are still made about the memory layout of wavefunctions, which for spin 1/2 and spin 1 particles are complex 6-dimensional arrays as in the original HELAS bib:helas and ALOHA bib:aloha implementations, from which CUDACPP is derived. This detail is relevant to the kernel splitting for different Feynman diagrams, as discussed in Sec. 3. no knowledge of the memory layouts used in the multi-event data buffers. These functions also ignore whether the calculations are performed in single or double precision, as I encapsulated this choice in the fptype type definition. Even more, these functions ignore whether the arithmetic operations within them are applied to scalar values on GPUs or CPUs or to SIMD vectors on CPUs: in the latter case, this is possible because simple operators like “+” are automatically understood by the compiler as vector operations when applied to CVE vectors of fptype, and are also implemented using SIMD CVE operations in their definition for vectors of custom types, notably the vectors of cxtype complex types. This design is extremely powerful because it has made it possible to use formally the same exact lines of software in the code-generated helicity amplitude functions like VVV1P0_1 or FFV1_0. I note in passing, however, that the functions computing wavefunctions for initial and final state particles, like IXXXXX, OXXXXX and VXXXXX, or their variations for massless or beam-collinear particles, are not code-generated from a model Lagrangian, but are instead hardcoded and often required particular care and successive iterations. Two complications that I had to address for SIMD code, in particular, are the following [^2] In this context, I find it important to stress that, in my experience, achieving a robust and performant CUDACPP implementation for SIMD CPUs was far from a trivial task, and was, in fact, much more complex than achieving it for GPUs. The difficulties for implementing IXXXXX-like functions are just two examples, and I already mentioned the fact that there are very strict constraints in the memory layouts for CPU SIMD, unlike those for GPUs. Another example is the implementation of mixed floating point mode, where I added a complex mechanism to merge/split two SIMD vectors of double into/from one SIMD vector of float at the boundary between dual amplitude and color sum calculations, in order to use the widest possible SIMD vectors in both cases and maximize efficiency. More generally, in my opinion, it is fair to say that achieving SIMD speedups through vectorization is always much more complicated than achieving speedups on GPUs. Code acceleration on GPUs, in fact, can be achieved even without lockstep processing (at the cost of some inefficiency from warp divergence), or without optimized memory layouts (at the cost of some inefficiency from the lack of coalesced data access), simply because a GPU has thousands of threads. One good feature of CUDACPP is that, by focusing on achieving C++ vectorization on SIMD CPUs, this has also improved the efficiency of the code on GPUs. : first, unlike VVV1P0_1 or FFV1_0, the IXXXXX-like functions include some if/else branching, which I reimplemented using vector masks in the SIMD case in order to fully exploit data parallel speedups (even if, from a performance point of view, this is only important for simple processes with few Feynman diagrams); second, as I performed all my functional testing with Floating Point Exception traps (FPEs) enabled in order to develop more robust code, I came across some crashes caused by the interplay of compiler optimizations, SIMD CVEs and FPE traps, which I addressed using volatile keywords and which pushed me to develop a large set of functional tests specifically for these functions.
@@
On top of the separation of the three software concerns above (data allocation, data access and arithmetic calculations), another important part of the design of the CUDACPP software was the implementation of kernel launching on GPUs and event loops on CPUs. My aim here was to keep as much as possible of the software logic and of the actual code identical for GPUs and CPUs, in order to simplify the iterative addition of new features and of bug fixes for both cases. In short, I addressed this by keeping also for SIMD CPUs the idea of a “grid” of events that are processed in one given iteration, and by explicitly subdividing this grid into SIMD event vectors and adding an explicit loop over them. Much more practically, and more importantly for the kernel splitting work described in the next section, the link between memory allocations and the execution of the ME computational engines is provided by two different incarnations of a MatrixElementKernel (MEK) class, one for GPUs (Device) and one for CPUs (Host). A singleton instance of the appropriate MEK class is constructed and used inside check.exe for standalone tests, or inside the Bridge component that is linked with Fortran in madevent for the full MG5aMC workflow. As shown in Table 1, it is the MatrixElementKernelDevice class that internally launches the monolithic sigmaKin kernel in the current CUDACPP. This is the starting point for the enhancements presented in the following section.
Process Feynman diagrams SDE channels Leading colors Color matrix Helicities Wave functions LOC (CPPProcess.cc)
$gg\!\rightarrow\!t\bar{t}$ 3 3 2 2 x 2 16 5 1424
$gg\!\rightarrow\!t\bar{t}g$ 16 15 6 6 x 6 32 12 1641
$gg\!\rightarrow\!t\bar{t}gg$ 123 105 24 24 x 24 64 26 3569
$gg\!\rightarrow\!t\bar{t}ggg$ 1240 945 120 120 x 120 128 121 31140
Table 2: Comparison of the computational complexity of the four physics processes considered in the tests described in this paper. The columns represent the following: (1) the physics process; (2) the number of Feynman diagrams; (3) the number of distinct SDE channels, i.e. the number of diagrams used in the MadEvent single diagram enhancement algorithm; (4) the number of leading QCD color flows; (5) the size of the color matrix; (6) the number of combinations of helicities for initial and final state particles (for the four processes considered, the number of good helicities contributing to the ME calculation coincides with the total number of helicities); (7) the number of wavefunctions to be computed for all external (initial and final state) and internal (propagator) particles; (8) the number of lines of code in the CUDAPP generated file CPPProcess.cc, in the ihel3 version of the software. All of these numbers are needed to describe the complexity of the ME calculation in a single madevent (or standalone) application, which is the main focus of this paper; the number of SDE channels, in addition, is one of the factors that determines how many instances of a madevent application must be launched in a full event generation workflow. The number of helicities determines how many GPU streams are used (in the ihel1 software). The number of leading colors determines the number of dual amplitudes that are computed using Feynman diagrams and used as input to the color sum; as of the ihel2 software, these are stored into and retrieved from GPU global memory. The size of the color matrix determines the complexity of the color sum calculation (addressed with a separate GPU kernel in the ihel2 software, and optionally via BLAS in the ihel3 software). The number of wavefunctions is particularly relevant to the ihel4 software, as these are stored into and retrieved from GPU global memory in that case.
3 New developments: GPU kernel splitting
@@
Since the ramp-up of ME throughputs with increasing GPU grid sizes is one of the main issues that this new work aims at addressing, or in any case a very good test of the effectiveness of these code changes, the results that I present in this section will essentially all consist of plots of that sort. One limitation of this work is that, due to lack of time, I will only show plots for a single CPU process, and only for the standalone application check.exe. This is unfortunate, because the practical benefits of any progress in this area would mainly be for the full MG5aMC workflow, using several madevent applications accessing the GPU in parallel. In the ACAT2022 presentation bib:acat2022 (see Fig. 1 therein), I had also studied the ramp-up of throughput as a function of the GPU grid size when several check.exe applications are launched in parallel, which seemed to indicate that the use of the GPU is more efficient in that case. Those results had been obtained using software containers prepared for the HEP-SCORE benchmarking project bib:bmk ; similar tests may be repeated in the future, when updated HEP-SCORE containers are preppared using more recent versions of CUDACPP, possibly including those presented in this paper.
Concerning my software development process, I note that for this new research I used the same methodology that I have been following for all of my work on CUDACPP during the last six years. In particular, I always prototyped, tested, fixed and optimized my changes using as baseline the code-generated software for one specific physics process, typically $gg\!\rightarrow\!t\bar{t}$ or $gg\!\rightarrow\!t\bar{t}gg$. In the case of this work on kernel splitting, I focused on the CUDA implementation but also tested the SIMD C++ version. I then committed all changes for the given process to git, until these reached a state that I considered reasonably complete. At that point, I backported my changes to the code-generating Python framework, and regenerated all physics processes. I then performed larger-scale functional and performance tests, also using different hardware implementations like AMD GPUs, and iterated until completion. This development process has been possible only because, while the git repository of the MG5aMC project bib:mg5amc-github mainly contains the hardcoded components of the framework and its code-generating engine, the git repository of madgraph4gpu project bib:mg4gpu-github contains not only the CUDACPP plugin but also the code-generated software for several physics processes. In my experience, this has been a key ingredient in the development of these latest enhancements to CUDACPP, but also much more generally of the whole plugin. In particular, I stress that I would never have been able to design and implement such large code changes directly in the code-generating Python code.
@@
Figure 3: Schematic representation of the CUDACPP ME engine (sigmaKin), in the last kernel splitting scenario described in this paper: (ihel4) Feynman diagrams as individual kernels.
Figure 4: Throughputs (ME/s) as a function of grid size for an NVidia V100 GPU (node itscrd90 at CERN). Higher is better. The 12 plots correspond to 4 physics processes in 3 floating point precisions. The number of threads per block is fixed to 32 (NVidia GPU warp size); the grid size is varied by changing the number of blocks. Each plot compares the different scenarios considered in this paper: (ihel0) current version before kernel splitting; (ihel1) helicity streams; (ihel2) color sum kernel; (ihel3) color sum kernel via host dispatcher; (ihel3b) color sum on cuBLAS, without TF32 math mode; (ihel4) Feynman diagrams as individual kernels.
Figure 5: Throughputs (ME/s) as a function of grid size for an AMD MI200 GPU (at LUMI). Higher is better. The 9 plots correspond to 3 physics processes in 3 floating point precisions. The number of threads per block is fixed to 256; the grid size is varied by changing the number of blocks. Each plot compares the different scenarios considered in this paper: (ihel0) current version before kernel splitting; (ihel1) helicity streams; (ihel2) color sum kernel; (ihel3) color sum kernel via host dispatcher; (ihel3b) color sum on hipBLAS via host dispatcher; (ihel4) Feynman diagrams as individual kernels.
@@
In the following subsections, I will describe the four sets of kernel splitting changes I developed: (ihel1) helicity streams; (ihel2) color sum splitting into separate GPU kernels; (ihel3) color sum host refactoring with optional BLAS offloading; and (ihel4) Feynman diagram splitting into separate GPU kernels. These developments are sequential, i.e. ihel4 includes ihel3, which includes ihel2, which includes ihel1. The software architecture of the sigmaKin ME engine is represented in Fig. 2 for the first three scenarios (where it is also compared to the situation before any kernel splitting, hereafter “ihel0”) and in Fig. 3 for the ihel4 case. Note that I will not provide new pseudo-code listings for my four kernel splitting developments, but in some cases I will refer to the pseudo-code for the current version ihel0 in Table 1 to point out what I changed or which technical issues I had to address. The results of my tests for these different versions of the software, which I will discuss more in detail in each subsection, are given in Fig. 4 for an NVidia V100 GPU at CERN and in Fig. 5 for an AMD MI200 GPU at LUMI, for three floating point precisions (double, mixed, float) and for four physics processes of increasing complexity ($g\!\rightarrow\!t\bar{t}, gg$$\rightarrow\!t\bar{t}g, gg$$\rightarrow\!t\bar{t}gg, gg$$g\!\rightarrow\!t\bar{t}ggg$), whose relevant parameters are described in detail in Table 2.

### v2

@@ 2.2 A closer look at the internals of the ME engine
    Arithmetic operations and other computational functions. Finally, the decoupling of data access from actual calculations has probably been one of the most important aspects of the whole CUDACPP software design. Just like the memory allocation classes, also the helicity amplitude functions like VXXXXX, VVV1P0_1 or FFV1_0 have almost [^2] One notable exception is that some assumptions are still made about the memory layout of wavefunctions, which for spin 1/2 and spin 1 particles are complex 6-dimensional arrays as in the original HELAS bib:helas and ALOHA bib:aloha implementations, from which CUDACPP is derived. This detail is relevant to the kernel splitting for different Feynman diagrams, as discussed in Sec. 4. no knowledge of the memory layouts used in the multi-event data buffers. These functions also ignore whether the calculations are performed in single or double precision, as I encapsulated this choice in the fptype type definition. Even more, these functions ignore whether the arithmetic operations within them are applied to scalar values on GPUs or CPUs or to SIMD vectors on CPUs: in the latter case, this is possible because simple operators like “+” are automatically understood by the compiler as vector operations when applied to CVE vectors of fptype, and are also implemented using SIMD CVE operations in their definition for vectors of custom types, notably the vectors of cxtype complex types. This design is extremely powerful because it has made it possible to use formally the same exact lines of software in the code-generated helicity amplitude functions like VVV1P0_1 or FFV1_0. I note in passing, however, that the functions computing wavefunctions for initial and final state particles, like IXXXXX, OXXXXX and VXXXXX, or their variations for massless or beam-collinear particles, are not code-generated from a model Lagrangian, but are instead hardcoded and often required particular care and successive iterations. Two complications that I had to address for SIMD code, in particular, are the following [^3] In this context, I find it important to stress that, in my experience, achieving a robust and performant CUDACPP implementation for SIMD CPUs was far from a trivial task, and was, in fact, much more complex than achieving it for GPUs. The difficulties for implementing IXXXXX-like functions are just two examples. I also already mentioned the fact that there are very strict constraints in the memory layouts for CPU SIMD, unlike those for GPUs; as described in Ref. bib:vchep2021 , in particular, this also implied the development of wrappers and adapters for vectors cxtype_v of complex numbers, as their real and imaginary parts must be stored as an SOA using two contiguous fptype_v arrays (RRRRIIII), rather than as an AOS (RIRIRIRI). Another example is the implementation of mixed floating point mode, where I added a complex mechanism to merge/split two SIMD vectors of double into/from one SIMD vector of float at the boundary between dual amplitude and color sum calculations, in order to use the widest possible SIMD vectors in both cases and maximize efficiency. More generally, in my opinion, it is fair to say that achieving SIMD speedups through vectorization is always much more complicated than achieving speedups on GPUs. Code acceleration on GPUs, in fact, can be achieved even without lockstep processing (at the cost of some inefficiency from warp divergence), or without optimized memory layouts (at the cost of some inefficiency from the lack of coalesced data access), simply because a GPU has thousands of threads. One good feature of CUDACPP is that, by focusing on achieving C++ vectorization on SIMD CPUs, this has also improved the efficiency of the code on GPUs. : first, unlike VVV1P0_1 or FFV1_0, the IXXXXX-like functions include some if/else branching, which I reimplemented using vector masks in the SIMD case in order to fully exploit data parallel speedups (even if, from a performance point of view, this is only important for simple processes with few Feynman diagrams); second, as I performed all my functional testing with Floating Point Exception traps (FPEs) enabled in order to develop more robust code, I came across some crashes caused by the interplay of compiler optimizations, SIMD CVEs and FPE traps, which I addressed using volatile keywords and which pushed me to develop a large set of functional tests specifically for these functions.
@@
On top of the separation of the three software concerns above (data allocation, data access and arithmetic calculations), another important part of the design of the CUDACPP software was the implementation of kernel launching on GPUs and event loops on CPUs. My aim here was to keep as much as possible of the software logic and of the actual code identical for GPUs and CPUs, in order to simplify the iterative addition of new features and of bug fixes for both cases. In short, I addressed this by keeping also for SIMD CPUs the idea of a “grid” of events that are processed in one given iteration, and by explicitly subdividing this grid into SIMD event vectors and adding an explicit loop over them. Much more practically, and more importantly for the kernel splitting work described in the next sections, the link between memory allocations and the execution of the ME computational engines is provided by two different incarnations of a MatrixElementKernel (MEK) class, one for GPUs (Device) and one for CPUs (Host). A singleton instance of the appropriate MEK class is constructed and used inside check.exe for standalone tests, or inside the Bridge component that is linked with Fortran in madevent for the full MG5aMC workflow. As shown in Table 1, it is the MatrixElementKernelDevice class that internally launches the monolithic sigmaKin kernel in the current CUDACPP. This is the starting point for the enhancements presented in the following sections.
2.3 A brief note on SIMD build modes
As mentioned previously, the new developments that I present in this paper are mainly optimizations that target GPUs specifically, but they are also relevant and useful for the CUDACPP implementation on vector CPUs. In this paper, in particular, I also present results from the five different SIMD build modes of CUDACPP, namely none, sse4, avx2, 512y and 512z. While these have already been described to some extent in earlier publications from our team bib:vchep2021 ; bib:zw2025 , I find it useful to take this paper as an opportunity to give a few additional and previously undocumented details.
The SIMD modes of CUDACPP were initially introduced targeting Intel CPUs. My design strategy was to define separate build modes, which on every specific system would allow not only the fastest calculation compatible with the hardware specs, but also an evaluation of the speedup achieved through the SIMD implementation of CUDACPP in vectorised C++. Largely speaking: none does not use any SIMD extensions; sse4 uses SSE4.2 instructions and 128-bit xmm registers; avx2 uses AVX2 instructions and 256-bit ymm registers; 512y uses AVX512 instructions and 256-bit ymm registers; 512z uses AVX512 instructions and 512-bit zmm registers. The 512y mode may appear surprising: the rationale behind it is, quite simply, that it is around 10% faster than avx2 on CPUs with the AVX512 instruction set, but it is also faster than 512z on AVX512 CPUs with a single FMA unit (such as Intel Xeon Silver 4216 CPUs); conversely, it is slower than 512z on AVX512 CPUs with two FMA unit (such as Intel Xeon Gold 6326 CPUs). This was first reported for the CUDACPP software in Ref. bib:ichep2022 and has been confirmed in many tests since then.
More in detail, each build mode corresponds to a well-defined choice of two parameters, the CPU instruction set, and the length of CUDACPP SIMD vectors:
  1. 1.
    The CPU instruction set is configured via the -march flag of the compiler, in the cudacpp.mk makefile. First and foremost, this must be well-defined because an application compiled for a given instruction set would crash if it is executed on a hardware that does not support it (for instance, if it is compiled for AVX512 but executed on a host that does not support it). While many different levels and sublevels of vectorization have been defined over time, such as AVX, AVX2, SSE4.1 or SSE4.2, the choice I made was to only support a very limited subset of these in CUDACPP. In particular: none uses -march=x86-64, which disables all SIMD extensions; sse4 uses -march=nehalem, which supports SSE4.2 extensions with 128-bit xmm registers; avx2 uses -march=haswell, which supports AVX2 extensions with 256-bit ymm registers; 512y and 512z both use -march=skylake-avx512, which supports a specific set of AVX512 extensions with 512-bit zmm registers. In this context, I chose not to foresee a build mode based on -march=native, which would offer the greatest set of hardware features available on a CPU, but is somewhat poorly defined. This is similar to what is done in some LHC experiments: LHCb, for instance, uses well-defined software architecture builds based on x86_64 Micro-Architecture Feature Levels bib:phoronix such as x86_64-v2. In retrospective, indeed a better choice would have been to configure -march based on x86_64 Feature Levels, but these were not yet widely used when I developed the SIMD implementation of CUDACPP in 2020.
  2. 2.
    The SIMD vector length is the core of the CUDACPP SIMD implementation, where it is encapsulated by a preprocessor variable MGONGPU_CPPSIMD and an equivalent C++ variable neppV. As described more in detail in Ref. bib:vchep2021 , in fact, the basic strategy of the CUDACPP SIMD implementation is the definition, using compiler vector extensions, of vectors fptype_v of neppV floating point variables of type fptype. The five build modes are defined as follows: none uses neppV=1, sse4 uses neppV=2 and neppV=4 in double and single precision, respectively; avx2 and 512y both use neppV=4 and neppV=8 in double and single precision; 512z uses neppV=8 and neppV=16 in double and single precision.
The SIMD instruction set and vector length in the five CUDACPP build modes are, of course, tightly related. The choice of the appropriate MGONGPU_CPPSIMD and neppV depends on the compiler options defined in the cudacpp.mk makefile. In most cases, notably none, sse4 and avx2, this depends on the preprocessor variables defined by the specific -march choice. In the case of 512y and 512z, however, which use the same -march, an additional compiler flag -DMGONGPU_PVW512 informs CUDACPP that it should “prefer vector width 512” in the 512z mode, i.e. zmm registers. In addition, the -mprefer-vector-width=256 flag is passed in the 512y mode, to enforce the use of ymm registers alone, although this is probably unnecessary.
Finally, I note that the names of the CUDACPP SIMD build modes are presently used also on non-Intel CPUs. For instance, on ARM CPUs the sse4 mode uses NEON vector instructions and 128-bit registers; this is defined via preprocessor directives that are currently being reviewed. This simplifies the user interfaces of the software, but is clearly a language abuse: eventually, it may be appropriate to switch to a better terminology. Apart from a rationalization of build modes, in any case, the port to new architectures should not pose any big challenge. Some opportunities for larger SIMD speedups may arise if new hardware architectures with larger vector registers became available, although this seems unlikely at the moment. From the CUDACPP point of view, this would only involve the definition of new build modes [^4] As a proof of concept, I already prepared an implementation with the values of neppV that would be appropriate to 1024-bit and 2048-bit registers, and tested its functionality. This is currently of no practical utility, also because, not surprisingly, it is slower than the existing CUDACPP SIMD modes. with larger values of neppV, but the overall implementation based on compiler vector extensions would remain essentially the same.
3 Kernel splitting I: helicity streams, color sum
@@
Physics process $g\!\rightarrow\!t\bar{t} gg$$\rightarrow\!t\bar{t}g gg$$\rightarrow\!t\bar{t}gg gg$$\rightarrow\!t\bar{t}ggg gg$$\rightarrow\!t\bar{t}gggg gg$$g\!\rightarrow\!t\bar{t}ggggg$
Feynman diagrams 3 16 123 1240 15495 231280
SDE channels 3 15 105 945 N/A N/A
Leading colors 2 6 24 120 720 5040
Color matrix 2 x 2 6 x 6 24 x 24 120 x 120 720 x 720 5040 x 5040
Helicities 16 32 64 128 256 512
Wave functions 5 12 26 121 750 5834
Code generation codebase ihel6p2 ihel6p2 ihel6p2 ihel6p2 ihel6p2 ihel6p1
Code generation mode mad mad mad mad sa sa
Code generation d.p.g. any any any any $\leq$100 $\leq$1000
Code generation d.p.f. any any any any $\geq$100 $\geq$1000
Code generation time 2s 2s 4s 32s 13m 30h
Feynman diagram groups 1 1 1 1 155 232
Feynman diagram files 1 1 1 1 155 232
LOC (color_sum.cc) 438 442 460 556 1156 5.4k
LOC (diagramsXYZ.cc) 135 307 1986 27k 652k 18.7M
Size (color_sum.cc) 24 kB 24 kB 28 kB 92 kB 2.6 MB 137 MB
Size (diagramsXYZ.cc) 8 kB 16 kB 84 kB 1.0 MB 28.2 MB 721 MB
Build time (five C++ modes) 6s 6s 6s 23s 3m 55m
Build time (CUDA DCDIAG=0) 11s 11s 14s 6m 9m failed
Build time (CUDA DCDIAG=1) 13s 13s 15s 3m 6m failed
Table 2: Comparison of the computational complexity of the six physics processes considered in the tests described in this paper. The rows represent the following: (1) the physics process; (2) the number of Feynman diagrams; (3) the number of distinct SDE channels, i.e. of the diagrams used in the MadEvent single diagram enhancement algorithm; (4) the number of leading QCD color flows; (5) the size of the color matrix; (6) the number of combinations of helicities for initial and final state particles (for these six processes, this coincides with the number of good helicities with non-zero contributions to MEs); (7) the number of wavefunctions to be computed for all external (initial and final state) and internal (propagator) particles; (8) the codebase used for code generation; (9) the code generation mode, “mad” for full generation including Fortran madevent and “sa” for standalone only; (9-10) the maximum number of diagrams per group (dpg) and the minimum number of diagrams per file (dpf) during code generation; (12) the code generation time on a typical CPU; (13-14) the resulting numbers of Feynman diagram groups and source code files; (15-17) the total number of lines of code and total size in the CUDAPP generated color_sum.cc and the sum of all diagramsXYZ.cc files; (19-21) the time for a parallel (-j) build of the software with a cold cache (CCACHE_RECACHE=1), either for all five C++ modes on an Intel Xeon Gold 6326, or for CUDA in each of two DCDIAG modes on an AMD EPYC 7313. All of these numbers are needed to describe the complexity of the ME calculation in a single standalone (or madevent) application, which is the main focus of this paper; the number of SDE channels, in addition, is one of the factors that determines how many instances of a madevent application must be launched in a full event generation workflow. The number of helicities determines how many GPU streams are used, as of the ihel1 software. The number of leading colors determines the number of dual amplitudes that are computed using Feynman diagrams and used as input to the color sum; as of ihel2, these are stored into and retrieved from GPU global memory. The size of the color matrix determines the complexity of the color sum calculation (addressed with a separate GPU kernel as of ihel2, and optionally via BLAS as of ihel3). The number of wavefunctions is particularly relevant to the ihel4-ihel6 software versions, which handle their storage and retrieval in GPU registers and/or GPU global memory in subtly different ways.
@@
Since the ramp-up of ME throughputs with increasing GPU grid sizes is one of the main issues that this new work aims at addressing, or in any case a very good test of the effectiveness of these code changes, the results that I present in this paper will mainly consist of plots of that sort. One limitation of this work is that, due to lack of time, I will only show plots for a single CPU process, and only for the standalone application check.exe. This is unfortunate, because the practical benefits of any progress in this area would mainly be for the full MG5aMC workflow, using several madevent applications accessing the GPU in parallel. In the ACAT2022 presentation bib:acat2022 (see Fig. 1 therein), I had also studied the ramp-up of throughput as a function of the GPU grid size when several check.exe applications are launched in parallel, which seemed to indicate that the use of the GPU is more efficient in that case. Those results had been obtained using software containers prepared for the HEP-SCORE benchmarking project bib:bmk ; it would be useful to repeat similar tests in the future, when updated HEP-SCORE containers are prepared using more recent versions of CUDACPP, possibly including those presented in this paper.
Concerning my software development process, I note that for this new research I used the same methodology that I have been following for all of my work on MG5aMC during the last four years, since the completion bib:ichep2022 of a fully functional CUDAPP code-generating plugin in the “epochX” development workflow at the end of 2021. In particular, I always prototyped, tested, fixed and optimized my changes using as baseline the code-generated software for one specific physics process, typically $gg\!\rightarrow\!t\bar{t}$ or $gg\!\rightarrow\!t\bar{t}gg$. In the case of this work on kernel splitting, I focused on the CUDA implementation but also tested the SIMD C++ version. I then committed all changes for the given process to git, until these reached a state that I considered reasonably complete. At that point, I backported my changes to the code-generating Python framework, and regenerated all physics processes. I then performed larger-scale functional and performance tests, also using different hardware implementations like AMD GPUs, and iterated until completion. This development process has been possible only because, while the git repository of MG5aMC bib:mg5amc-github mainly contains the hardcoded components of the framework and its code-generating engine, the git repository of the madgraph4gpu project bib:mg4gpu-github contains not only the CUDACPP plugin but also the code-generated software for several physics processes. In my experience, this has been a key ingredient in the development of these latest enhancements to CUDACPP, but also much more generally of the whole plugin. In particular, I stress that I would never have been able to design and implement such large code changes directly in the code-generating Python code.
@@
In the following subsections, I will describe the first batch of three sets of kernel splitting changes that I developed, and which have already been integrated in a new production release of CUDACPP: (ihel1) helicity streams; (ihel2) color sum splitting into separate GPU kernels; (ihel3) color sum host refactoring with optional BLAS offloading. As shown in Fig. 1, these developments are sequential, i.e. ihel3 includes ihel2, which includes ihel1. The software architecture of the sigmaKin ME engine for these three scenarios is represented in Fig. 3, where it is also compared to the situation before any kernel splitting, ihel0. Note that I will not provide new pseudo-code listings for these kernel splitting developments, but in some cases I will refer to the pseudo-code for version ihel0 in Table 1 to point out what I changed or which technical issues I had to address. The results of my tests for these different versions of the software, as well as for those in my second batch of changes described in the next section, are given in Fig. 4 for an NVidia V100 GPU at CERN, in Fig. 5 for an AMD MI200 GPU at LUMI, and in Fig. 6 for an Intel Xeon Silver 4216 CPU, for three floating point precisions (double, mixed, float) and for four physics processes of increasing complexity ($g\!\rightarrow\!t\bar{t}, gg$$\rightarrow\!t\bar{t}g, gg$$\rightarrow\!t\bar{t}gg, gg$$g\!\rightarrow\!t\bar{t}ggg$), whose relevant parameters are described in detail in Table 2.
Figure 4: Throughputs (ME/s) as a function of grid size for an NVidia V100 GPU (node itscrd90 at CERN). Higher is better. The 12 plots correspond to 4 physics processes in 3 floating point precisions. The number of threads per block is fixed to 32 (NVidia GPU warp size); the grid size is varied by changing the number of blocks. Each plot compares different scenarios considered in this paper: (ihel0) release v1.00.02, before kernel splitting; (ihel1) helicity streams; (ihel3) color sum kernel; (ihel3b) cuBLAS color sum; (ihel3p1) release v1.01.01, color sum kernel; (ihel3b) release v1.01.01, cuBLAS color sum; (ihel4) Feynman diagrams as individual kernels; (ihel6p2) diagram groups, color sum kernel; (ihel6p2b) diagram groups, cuBLAS color sum. For ihel6p2 and ihel6p2b, all four processes were generated with a single diagram group, which was executed as a kernel at runtime, without graphs.
Figure 5: Throughputs (ME/s) as a function of grid size for an AMD MI200 GPU (at LUMI). Higher is better. The 9 plots correspond to 3 physics processes in 3 floating point precisions. The number of threads per block is fixed to 256; the grid size is varied by changing the number of blocks. Each plot compares different scenarios considered in this paper: (ihel0) release v1.00.02, before kernel splitting; (ihel1) helicity streams; (ihel3) color sum kernel; (ihel3b) hipBLAS color sum; (ihel3p1) release v1.01.01, color sum kernel; (ihel3b) release v1.01.01, hipBLAS color sum; (ihel4) Feynman diagrams as individual kernels; (ihel6p2) diagram groups, color sum kernel; (ihel6p2b) diagram groups, hipBLAS color sum. For ihel6p2 and ihel6p2b, all four processes were generated with a single diagram group, which was executed as a kernel at runtime, without graphs.
Figure 6: Throughputs (ME/s) as a function of CUDACPP SIMD build mode for an Intel Xeon Silver 4216 CPU (node itscrd90 at CERN). Higher is better. The 12 plots correspond to 4 physics processes in 3 floating point precisions. Each plot compares different scenarios considered in this paper: (ihel0) release v1.00.02, before kernel splitting; (ihel3p1) release v1.01.01; (ihel4) Feynman diagrams as individual kernels; (ihel6p2) diagram groups, using code generated with a single diagram group.